# Wang Park Medennikov Wang Dhawan Huang Rao Koluguri Balam Ginsburg NVIDIAUSA

### v1

@@ 1 Introduction
In parallel, speaker-attributed ASR [19] has significantly improved ASR performance for overlapped multi-talker speech. Recently, alignment-free training (AFT) [20] has shown that high performance can be achieved without explicit alignment by training and predicting each speaker’s speech separately. These advancements mark substantial progress in multi-talker ASR.
@@
On the contrary, we leverage this mechanism in reverse to reinforce the propensity of adhering to a specific speaker. Our proposed technique achieves this by injecting a learnable speaker kernel into the pre-encode layer of the ASR encoder [21]. This enables the ASR encoder to detect speech presence and utilize the speech kernel to adapt the encoder states dynamically. As a result, the encoder becomes more responsive to the targeted speaker’s speech characteristics. However, this approach requires deploying one model instance per speaker, meaning the number of model instances must match the number of speakers. While this necessitates additional computational resources, it significantly enhances multi-talker ASR performance, achieving state-of-the-art error rates on benchmark datasets for multi-talker ASR.
@@ 2 Proposed Method
Typically, target-speaker ASR systems rely on target speaker embeddings or enrollment audio to extract and utilize speaker-specific information from multi-talker utterances. However, the performance of such systems is highly dependent on the quality of the provided queries. For instance, a clean and noise-free query audio is generally preferred to achieve optimal results in target-speaker ASR tasks. In this work, we introduce a novel SSA approach that enables the model to adapt to a specific speaker using only the corresponding speech activities, eliminating the need for high-quality external queries. This method leverages the inherent speaker characteristics present in the input audio, allowing the model to dynamically adjust its focus to better recognize the target speaker’s speech. By doing so, our approach reduces the dependency on external resources and enhances the robustness of the system in real-world scenarios where high-quality queries or speaker representations may not always be available.
@@ 2.1 Self-Speaker Adaptation
By leveraging the provided speech activities, the model dynamically adapts to the specific speaker, enabling robust recognition even in multi-speaker scenarios. Notably, the training process remains nearly identical to that of a conventional single-speaker ASR system, including the training objective and decoding procedure. The only distinctions lie in the model architecture, which includes the speaker injection module, and the use of multi-speaker speech data for training. This approach ensures compatibility with existing ASR frameworks while enhancing the model’s ability to handle target-speaker scenarios effectively.
@@ 2.2 Repurposing for Multi-Talker ASR
Figure 3 shows the t-SNE plot of the ASR encoder state (activation at the last layer) for each token, where 2mix-spk0 and 2mix-spk1 are injected with the first and second speaker’s kernel, respectively. It is important to note that these ASR encoder representations, 2mix-spk0 and 2mix-spk1, are derived from the same model and audio recording. We observe that the variability introduced by speaker injection is smaller than the distance between tokens; however, there is still a clear distinction between each speaker’s kernel, enabling the model to decode overlapping speech from two or three speakers.
@@ 3.1 Datasets and Evaluation Metrics
Figure 3: t-SNE plot of Fast Conformer’s last encoder states (embeddings) from an identical audio input (“there are the characters you can see”), using the same model weights but processed with different speaker kernels. The plot contrasts how ASR embeddings differ when generated using kernel $spk0$ (triangles) versus $spk1$ (circles), likely reflecting distinct speaker focuses.

### v2

@@ 1 Introduction
In parallel, speaker-attributed ASR [19] has significantly improved ASR performance for overlapped multi-talker speech. Recently, alignment-free training (AFT) [20] has shown that high performance can be achieved without explicit alignment by training and predicting each speaker's speech separately. These advancements mark substantial progress in multi-talker ASR. In addition, a more sophisticated but streamlined target-speaker-based approach, called Diarization-Conditioned Whisper (DiCoW) [21], was proposed, in which diarization results condition the ASR model through a query-key biasing scheme.
@@
On the contrary, we leverage this mechanism in reverse to reinforce the propensity of adhering to a specific speaker. Our proposed technique achieves this by injecting a learnable speaker kernel into the pre-encode layer of the ASR encoder [22]. This enables the ASR encoder to detect speech presence and utilize the speech kernel to adapt the encoder states dynamically. As a result, the encoder becomes more responsive to the targeted speaker's speech characteristics. However, this approach requires deploying one model instance per speaker, meaning the number of model instances must match the number of speakers. While this necessitates additional computational resources, it significantly enhances multi-talker ASR performance, achieving state-of-the-art error rates on benchmark datasets for multi-talker ASR.
@@ 2 Proposed Method
Typically, target-speaker ASR systems rely on target speaker embeddings or enrollment audio to extract and utilize speaker-specific information from multi-talker utterances. However, the performance of such systems is highly dependent on the quality of the provided queries. For instance, a clean and noise-free query audio is generally preferred to achieve optimal results in target-speaker ASR tasks. In this work, we introduce a novel SSA approach that enables the model to adapt to a specific speaker using only the corresponding speech activities, eliminating the need for high-quality external queries. This method leverages the inherent speaker characteristics present in the input audio, allowing the model to dynamically adjust its focus to better recognize the target speaker's speech. By doing so, our approach reduces the dependency on external resources and enhances the robustness of the system in real-world scenarios where high-quality queries or speaker representations may not always be available.
@@ 2.1 Self-Speaker Adaptation
By leveraging the provided speech activities, the model dynamically adapts to the specific speaker, enabling robust recognition even in multi-speaker scenarios. Notably, the training process remains nearly identical to that of a conventional single-speaker ASR system, including the training objective and decoding procedure. The only distinctions lie in the model architecture, which includes the speaker injection module, and the use of multi-speaker speech data for training. This approach ensures compatibility with existing ASR frameworks while enhancing the model's ability to handle target-speaker scenarios effectively.
@@ 2.2 Repurposing for Multi-Talker ASR
Figure 3 shows the t-SNE plot of the ASR encoder state (activation at the last layer) for each token, where 2mix-spk0 and 2mix-spk1 are injected with the first and second speaker's kernel, respectively. It is important to note that these ASR encoder representations, 2mix-spk0 and 2mix-spk1, are derived from the same model and audio recording. We observe that the variability introduced by speaker injection is smaller than the distance between tokens; however, there is still a clear distinction between each speaker's kernel, enabling the model to decode overlapping speech from two or three speakers.
@@ 3.1 Datasets and Evaluation Metrics
Figure 3: t-SNE plot of Fast Conformer's last encoder states (embeddings) from an identical audio input (``there are the characters you can see''), using the same model weights but processed with different speaker kernels. The plot contrasts how ASR embeddings differ when generated using kernel $spk0$ (triangles) versus $spk1$ (circles), likely reflecting distinct speaker focuses.